OpenStudy (love_333):

Help? Will award medal for answer and walk through of how to find the answer. The measure of the larger of two adjacent angles is 12 degrees more than three times the measure of the smaller angle. If the two angles are also supplementary, what is the measure of the smaller angle?

OpenStudy (love_333):

A. 42 degrees B. 56 degrees C. 64 degrees D. 138 degrees

OpenStudy (love_333):

@firekat97 Can you tell me how you got that?

OpenStudy (firekat97):

wait no I read it wrong lol sooo lets assume two angles; a and b so lets say a is the larger one. so in the first bit we understand that a = 12 + 3b and in the second part we are told that they are both supplementary (add up to 180) therefore a + b = 180 so now we have two simultaneous equations. (1) a = 12 + 3b (2) a + b = 180 ...... a = 180 -b therefore 180 - b = 12 + 3b (180 - 12) = (3b + b) 168 = 4b therefore b = 42 so the answer is actually 42 not 56 sorry